### v1

@@ Appendix A Additional figures
Figure 18: Hertzsprung-Russel diagram similar to Fig. 3 but including only the 216 stars from our Gaia-detected sample having TESS light curves. Each star is colour coded according to its light curve shape and content in the Fourier domain. The stars classified as other are discussed in the main text.
Figure 19: Same as Fig. 8 but based on the probabilistic assignment of mode identifications from multi-colour photometry for all modes with $p>0.5$.
@@
Figure 20: Hertzsprung-Russel diagram similar to Fig. 3 featuring only the stars with identified modes from multi-colour photometry and colour-coded by their mode degree: $l=0$ (green), $l=1$ (blue), $l=2$ (orange).
Appendix B Identification of the radial order
@@
For each star, we produced a diagnostic plot similar to Fig. 21. It shows for all selected five radial orders the positions of all twelve considered low order radial, p, and g modes over-plotted on the frequency spectrum. In these diagrams, we examined the positions of these modes with respect to all identified (and unidentified low-SNR) mode frequencies.
In the chosen example, the lowest order fit to the dominant mode ($n=-2$, orange), is able to approximately explain detected modes left to the main frequency. However, it fails to reproduce the highest detected frequency in Fig. 21. Both the $n=-1$ (blue) and $n=0$ (green) fit to the dominant frequency and other modes near that highest frequency. Yet, we see that the $g_{1}$-mode matches the modes near 4.5 d-1 better. We associate the $g_{1}$ ($l=1$) mode with the secondary peak next to the identified frequency. The modes shown do not include any modes near 3.9 d-1 but higher-order g-modes might be found there for the $g_{1}$ model. In the best fitting p-mode models (purple and brown) the predicted mode frequencies do not match the observed frequencies.
In conclusion, we identified the dominant mode in this exemplary case as a $g_{1}$-mode due to the better fit of the model’s mode spectrum to the overall population of visible modes in the frequency spectrum. As additional information, we obtained the $\chi^{2}$ for the fitted model based on $T_{\mathrm{eff}}$, $\log L$, and $f_{0}$. In this case, the $\chi^{2}$ is lowest for the $g_{1}$ model giving additional confirmation for the model preferred from visual inspection.
@@
Figure 21: Diagnostic plot to identify the radial order $n$. In the background, we show in black to frequency spectrum of TIC 435101777. The lines indicate the positions of extracted independent frequencies. The short coloured lines show the positions of the modes for the best fitting model of different radial orders. From bottom to top, we show we show $n=-2,-1,0,1,2$ in orange, blue, green, purple, brown, respectively. The horizontal grey line crossing the main peak, indicates the identified splitting.
Appendix C Properties of the modelling results
C.1 Recovered values
@@
By comparing the input and recovered values from the grid modelling, we can assess the accuracy and understand potential biases of the modelling. The three panels of Fig. 22 show the measured and recovered $\log T_{\mathrm{eff}}$, $\log L$, and $f_{0}$. There is a slight bias towards higher effective temperatures as the majority of best fitting modes are located above the line of unity. A similar effect can be found for the luminosity, where the lower luminosity stars tend to be fitted with higher luminosity models. However, all models are within the estimated uncertainty of the measurements. The frequencies are mostly well recovered and deviate typically less than the uncertainty interval (0.5 d-1) from the line of unity. We find four stars well outside of this region. These stars are among the stars with the largest differences in the luminosity and effective temperature as well. Hence, no optimal model could be found for these stars with the adopted input physics chosen for the grid.
@@
Figure 22: Comparison between the measured values and the properties recovered from the grid modelling. From left to right we show the effective temperature, the luminosity, and the pulsation frequency of the dominant mode. The dashed lines mark the line of unity and the dotted lines the uncertainty interval used in our fitting procedure. The colour code indicates the main pulsation mode: $l=0$ green, $l=1$ blue, $l=2$ orange.
C.2 Ages
@@
Figure 23: Comparison between ages for stars in open clusters and our asteroseismic ages. Members of NGC 637 are shown with square and these of NGC 5606 with triangles. The typical uncertainty is given in the upper left corner.

### v2

@@ Appendix B Additional Hertzsprung-Russell diagrams and mode identification
Figure 21: Hertzsprung-Russell diagram similar to Fig. 3 but including only the 216 stars from our Gaia-detected sample having TESS light curves. Each star is colour-coded according to its light curve shape and content in the Fourier domain with purple indicating mono-periodic stars, olive multi-periodic, and red hybrid pulsators. The stars classified as other are shown in cyan and are discussed in the main text.
@@
Figure 22: Hertzsprung-Russell diagram similar to Fig. 3 featuring only the stars with identified modes from multi-colour photometry and colour-coded by their mode degree: $l=0$ (green), $l=1$ (blue), $l=2$ (orange).
Figure 23: Same as Fig. 9 but based on the probabilistic assignment of mode identifications from multi-colour photometry for all modes with $p>0.6$.
Appendix C Identification of the radial order
@@
For each star, we produced a diagnostic plot similar to Fig. 24. It shows for all relevant radial order out of the selected five radial orders the positions of all twelve considered low order radial and zonal p- and g-modes over-plotted on the frequency spectrum. In these diagrams, we examined the positions of these modes with respect to all identified (and unidentified low-SNR) mode frequencies.
In the chosen example, the radial order $n=1$ is able to approximately explain detected modes at frequencies both higher and lower than the identified dominant mode. However, it fails to reproduce the highest detected frequency in Fig. 24. Hence, we can conclude that TIC 105756939 is pulsating with $(n,l)=(1,1)$.
@@
Figure 24: Diagnostic plot to identify the radial order $n$. In the background, we show in black the frequency spectrum of TIC 105756939. The thin red vertical lines indicate the positions of extracted frequencies. The short coloured lines show the positions of the model predictions for the modes of different radial orders. From bottom to top, we show we show $n=-2,-1,1,2$ in orange, blue, purple, brown, respectively.
Appendix D Properties of the modelling results
D.1 Recovered values
@@
By comparing the input and recovered values from the grid modelling, we can assess the accuracy and understand potential biases of the modelling. The three panels of Fig. 25 show the measured and the differences to the recovered $\log T_{\mathrm{eff}}$, $\log L$, and $f_{0}$. There is a slight bias towards higher effective temperatures as the majority of best fitting modes are located below zero. A similar effect can be found for the luminosity, where the lower luminosity stars tend to be fitted with higher luminosity models. We also note that stars close to the edges of the model grid are affected by the lack of models and their luminosity is shifted towards the model grid, leading to the structure seen at low and high luminosities. However, all models are within the estimated uncertainty of the measurements. The frequencies are well recovered and deviate less than the uncertainty interval (0.1 d-1) from the measurements.
@@
Figure 25: Comparison between the measured values and the properties recovered from the grid modelling. From top to bottom we show the effective temperature, the luminosity, and the pulsation frequency of the dominant mode. The dotted lines mark the uncertainty interval used in our fitting procedure. The colour code indicates the main pulsation mode: $l=0$ green, $l=1$ blue, $l=2$ orange.
D.2 Ages
@@
Figure 26: Comparison between ages for stars in open clusters and our asteroseismic ages. Members of NGC 637 are shown with squares and those of NGC 5606 with triangles. The typical uncertainty is given in the upper left corner.
Appendix E Light curves and frequency spectra
E.1 Light curves for all studied stars
Figure 27: Light curves (left) and periodograms (right) for all studied stars. The light curves show only the first available TESS sector for every star for illustrative purpose. The periodogram is based on all available TESS data. For each star, we indicate its light curve classification in the top right.
Figure 28: continued.
Figure 29: continued.
Figure 30: continued.
Figure 31: continued.
Figure 32: continued.
Figure 33: continued.
Figure 34: continued.
Figure 35: continued.
Figure 36: continued.
Figure 37: continued.
Figure 38: continued.
Figure 39: continued.
E.2 Stillstand stars
In this section, we show the remaining stars that display a stillstand in their dominant pulsation mode not yet shown in the main text. Four of these are displayed in Fig. 40.
Figure 40: Same as Fig. 10, showing an additional four stillstand stars not yet shown in the main text.
[…]
Figure 41: TIC 395218466 is an eccentric eclipsing binary with a dominant radial mode displaying a stillstand. In the upper left panel, the unbinned light curve is phase folded over its dominant pulsation period of 0.26433 d like in Fig. 10. In the upper right panel, the light curve is instead folded over the orbital period of 2.657983 d after all extracted signals but the orbital harmonics were removed. The light curve shown in this figure is not binned to a 30 minute cadence as the eclipses are more clearly defined in the unbinned light curve. The bottom panel shows its original Fourier spectrum.
E.3 Splittings
Here, we present all identified complete split multiplets of the pulsations modes of the studied $\beta$ Cep stars. As seen from in Sect. E.1, some stars show more such multiplets that will be discussed elsewhere.
Figure 42: Splittings identified in the sample of $\beta$ Cep stars and used in the multicolour photometry. The connected, red markers indicate the splitting and the red markers at the top show all extracted frequencies. Information on the stars can be found in the box in each panel.
Figure 43: continued.
Figure 44: Additional splittings found in the sample of $\beta$ Cep stars. The symbols are identical to Fig. 42.